Entomologists have discovered that a linear relationship exists between the rate of chirping of crickets of a certain species and the air temperature. When the temperature is 70°F, the crickets chirp at the rate of 193 chirps/min, and when the temperature is 90°F, they chirp at the rate of 293 chirps/min.

(a) Find an equation giving the relationship between the air temperature x and the number of chirps per min y of the crickets.
y =  

  

(b) Use this function to determine the rate at which the crickets chirp when the temperature is 99°F.
__________ chirps/min
Answer #1

@ Y= axtb 1. Put y = 193, X= 70 and y=893x = 90 » 193= 70a4b 293=90a4b 2993-193= 90 at b-709-b = a Loo 35 b= 193-350 = -157 H